How many moles of carbon dioxide are formed when 27.6g of potassium carbonate reacts? Calulate the mass of carbon dioxide formed


1
Expert's answer
2021-03-30T07:53:11-0400

K2CO3 "\to" K2O + CO2

moles of K2CO3 = 27.6/138 = 0.2 moles

so 0.2 moles of CO2 is formed

mass of CO2 formed = 0.2×44 = 8.8g